# 22email: {gonzalo.mancera, aythami.morales, julian.fierrez, ruben.tolosana, alejandro.penna, miguel.lopezd

### v1

@@ Abstract
The use of Natural Language Processing (NLP) in high-stakes AI-based applications has increased significantly in recent years, especially since the emergence of Large Language Models (LLMs). However, despite their strong performance, LLMs introduce important legal/ethical concerns, particularly regarding privacy, data protection, and transparency. Due to these concerns, this work explores the use of Named-Entity Recognition (NER) to facilitate the privacy-preserving training (or adaptation) of LLMs. We propose a framework that uses NER technologies to anonymize sensitive information in text data, such as personal identities or geographic locations. An evaluation of the proposed privacy-preserving learning framework was conducted to measure its impact on user privacy and system performance in a particular high-stakes and sensitive setup: AI-based resume scoring for recruitment processes. The study involved two language models (BERT and RoBERTa) and six anonymization algorithms (based on Presidio, FLAIR, BERT, and different versions of GPT) applied to a database of $24$,$000$ candidate profiles. The findings indicate that the proposed privacy preservation techniques effectively maintain system performance while playing a critical role in safeguarding candidate confidentiality, thus promoting trust in the experimented scenario. On top of the proposed privacy-preserving approach, we also experiment applying an existing approach that reduces the gender bias in LLMs, thus finally obtaining our proposed Privacy- and Bias-aware LLMs (PB-LLMs). Note that the proposed PB-LLMs have been evaluated in a particular setup (resume scoring), but are generally applicable to any other LLM-based AI application.
@@ 1 Introduction
Although offering operational advantages, these technologies also face challenges related to fairness [51, 11, 40] and privacy [7, 28, 61]. Text data often contains personally identifiable information (PII) [49] - names, addresses, and demographic details — which, if mishandled, could expose individuals or introduce bias into decision-making processes. Recent regulatory initiatives, such as the AI Act [^1] https://artificialintelligenceact.eu/ approved by the European Parliament in June 2024 [14] and the White House Memorandum on AI leadership [57], seek to address these risks by promoting transparency, accountability, and fairness in AI systems.
@@
This paper addresses this challenge by proposing a generalizable approach to anonymize input text while preserving system performance, in a similar way related works remove unnecessary sensitive data while maintaining good performance in a main AI task [34], in this case particularly applied to NLP tasks based on LLMs (e.g., document layout analysis [41] or document topic classification [42]). Our method uses NER to detect and remove sensitive entities, minimizing privacy risks. The main idea is depicted in Fig. 1, where we show two branches: with and without removing the sensitive entities found using Named-Entity Recognition. In addition to that privacy enhancement via NER, we also apply and experiment the bias reduction methods recently proposed in [40] to derive our final Privacy- and Bias-aware models (PB-LLMs).
@@ 5.2 Privacy- and Bias-aware Models
The results shown before have been evaluated for the task of Occupancy Classification (see Sect.4.3). We now evaluate our PB-LLM approach (incorporating both NER-based privacy enhancement and gender bias reduction, similar to Table2) for the other task studied in the present paper: Candidate Scoring. For this second task of candidate scoring, we follow the same protocol of [39, 40], generating a top 100 list of candidates based on the estimated candidate scores $\hat{y}_{i}$ (see the right output in Fig.2). Recall that we used the men-biased dataset from FairCVdb, so the top 100 list is expected to have more men than women for similar competencies given the gender-biased expert score annotations. Eliminating personal data and gender cues in biographies using our PB-LLMs, Table 3 shows that even with these gender-biased training data, one can generate privacy-preserving bias-free AI models (useful enough for NLP tasks like the two tasks studied here: Occupancy Classification and Candidate Scoring).

### v2

@@ Abstract
The use of Natural Language Processing (NLP) in high-stakes AI-based applications has increased significantly in recent years, especially since the emergence of Large Language Models (LLMs). However, despite their strong performance, LLMs introduce important legal/ethical concerns, particularly regarding privacy, data protection, and transparency. Due to these concerns, this work explores the use of Named-Entity Recognition (NER) to facilitate the privacy-preserving training (or adaptation) of LLMs. We propose a framework that uses NER technologies to anonymize sensitive information in text data, such as personal identities or geographic locations. An evaluation of the proposed privacy-preserving learning framework was conducted to measure its impact on user privacy and system performance in a particular high-stakes and sensitive setup: AI-based resume scoring for recruitment processes. The study involved two language models (BERT and RoBERTa) and six anonymization algorithms (based on Presidio, FLAIR, BERT, and different versions of GPT) applied to a database of $24$,$000$ candidate profiles. The findings indicate that the proposed privacy preservation techniques effectively maintain system performance while playing a critical role in safeguarding candidate confidentiality, thus promoting trust in the experimented scenario. On top of the proposed privacy-preserving approach, we also experiment applying an existing approach that reduces the gender bias in LLMs, thus finally obtaining our proposed Privacy- and Bias-aware LLMs (PBa-LLMs). Note that the proposed PBa-LLMs have been evaluated in a particular setup (resume scoring), but are generally applicable to any other LLM-based AI application.
@@ 1 Introduction
Although offering operational advantages, these technologies also face challenges related to fairness [51, 11, 40] and privacy [7, 28, 61]. Text data often contains personally identifiable information (PII) [49] - names, addresses, and demographic details that, if mishandled, could expose individuals or introduce bias into decision-making processes. Recent regulatory initiatives, such as the AI Act [^1] https://artificialintelligenceact.eu/ approved by the European Parliament in June 2024 [14] and the White House Memorandum on AI leadership [57], seek to address these risks by promoting transparency, accountability, and fairness in AI systems.
@@
This paper addresses this challenge by proposing a generalizable approach to anonymize input text while preserving system performance, in a similar way related works remove unnecessary sensitive data while maintaining good performance in a main AI task [34], in this case particularly applied to NLP tasks based on LLMs (e.g., document layout analysis [41] or document topic classification [42]). Our method uses NER to detect and remove sensitive entities, minimizing privacy risks. The main idea is depicted in Fig. 1, where we show two branches: with and without removing the sensitive entities found using Named-Entity Recognition. In addition to that privacy enhancement via NER, we also apply and experiment the bias reduction methods recently proposed in [40] to derive our final Privacy- and Bias-aware models (PBa-LLMs).
@@ 5.2 Privacy- and Bias-aware Models
The results shown before have been evaluated for the task of Occupancy Classification (see Sect.4.3). We now evaluate our PBa-LLM approach (incorporating both NER-based privacy enhancement and gender bias reduction, similar to Table2) for the other task studied in the present paper: Candidate Scoring. For this second task of candidate scoring, we follow the same protocol of [39, 40], generating a top 100 list of candidates based on the estimated candidate scores $\hat{y}_{i}$ (see the right output in Fig.2). Recall that we used the men-biased dataset from FairCVdb, so the top 100 list is expected to have more men than women for similar competencies given the gender-biased expert score annotations. Eliminating personal data and gender cues in biographies using our PBa-LLMs, Table 3 shows that even with these gender-biased training data, one can generate privacy-preserving bias-free AI models (useful enough for NLP tasks like the two tasks studied here: Occupancy Classification and Candidate Scoring).